OpenStudy (anonymous):

.Write the equation 16x + 11y = −88 in slope-intercept form.

OpenStudy (anonymous):

Do you know what the slope intercept form is?

OpenStudy (anonymous):

y = mx + b is the slope intercept form where m is slope & b is intercept we have 16x + 11y = -88 so first we move 16x to right hand side therfore 11y = -16x -88 now we move 11 to right hand side so we will get y = -16x/11 - 88/11 by comparing with above equation we get m = -16/11

OpenStudy (anonymous):

It's actually \[y=\frac{ -16 x}{ 11 } -8 \] in slope intercept form

OpenStudy (anonymous):

slope intercept form is \[y=mx+b\]